how do you solve 4(5x-2)=92


Answer by dfrazzetto(283)

4(5x-2)=92


distribute the 4:


20x - 8 = 92


add 8 to both sides:


20x = 100


divide by 20:


x = 5


CHECK:
4(5(5) - 2) = 92
4(25-2) = 92
4(23) = 92 √